OpenStudy (anonymous):

ill fan you and give you a medal if someone helps me(: Find the solution of this system of equations. Separate the x-and y- values with a comma. x=10+y 12x-9y=9

OpenStudy (anonymous):

@LunyMoony

OpenStudy (anonymous):

@ganeshie8 , danish071996 said that you might be able to help me

OpenStudy (math&ing001):

Hi, we'll need to plug in "x=10+y" in the second equation and solve for y, can you do that ?

OpenStudy (anonymous):

how would you do that?

OpenStudy (math&ing001):

Simply by replacing "x" with its value "10+y" like this : 12(10+y) - 9y = 9 Can you solve for y now ?

OpenStudy (anonymous):

I can try give me a min

OpenStudy (math&ing001):

Sure, you can write it on paper and take a picture of it if it's easier for you.

OpenStudy (anonymous):

is there a simpler way to find y besides putting in random a_ss seeing if they work

OpenStudy (anonymous):

y=-3

OpenStudy (anonymous):

actually its 3 not (-3)

OpenStudy (math&ing001):

How did you get that ?

OpenStudy (anonymous):

just put in random numbers from 1-10 positive and negative and -s - so I made the 3 a + and it worked

OpenStudy (math&ing001):

Plugging in random numbers is not how you do it. it's gonna cost you a lot of time and you're not sure to get a correct answer. Here's how you do it : 12(10+y) - 9y = 9 120 + 12y - 9y = 9 (using FOIL) -120 -120 12y - 9y = 9 - 120 3y = -111 y = -111/3 y = -37

OpenStudy (anonymous):

im not good in math lol but can you help me on like 6 more?

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

Well this one is not over yet, we still have to find x.

OpenStudy (anonymous):

ugh! ;p okay

OpenStudy (math&ing001):

Ok, plug in y=-37 in the first equation x=10+y, and solve for x. Just like we did before.

OpenStudy (math&ing001):

And this one you'll need to find for yourself :) I'll walk you through if you need any help.

OpenStudy (math&ing001):

@lexikaylynn

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x=-27

OpenStudy (math&ing001):

Yep, correct !

OpenStudy (anonymous):

so the answer is -27, -37??

OpenStudy (math&ing001):

Yes (x,y) = (-27,-37)

OpenStudy (anonymous):

thanks (: do you think you can help me on like 6 more?

OpenStudy (math&ing001):

I have to go now, but I'll help you as soon as I get back.

OpenStudy (anonymous):

okay when will that be?